OpenStudy (anonymous):

I'm doing parent functions for calculus. What are the domains and ranges for: 1. f(x)=a^x 2. f(x)=logaX 3. f(x)=sinX 4. f(x)=cosX 5. f(x)=tanX

OpenStudy (anonymous):

1) all real numbers. you can raise a to any power 2) \[(0,\infty)\] you cannot take the log of a negative number because \[a^x=\text{something negative}\] has no solution

OpenStudy (anonymous):

3) sine all real number

OpenStudy (anonymous):

4) cosine all real numbers

OpenStudy (anonymous):

5) tangent all numbers except those for which \[\cos(x)=0\] those would be numbers like \[\frac{\pi}{2},\frac{3\pi}{2},\frac{5\pi}{2}\] i.e. all real numbers except \[x=\frac{(2k-1)\pi}{2}\] for integer k

OpenStudy (anonymous):

Thanks a lot. but for f(x)=logaX , what would the graph look like.